How many years will it take an investment of $1000 to double itself when interest is compunded annually at 6%?
Answer by stanbon(75887)

How many years will it take an investment of $1000 to double itself when interest is compunded annually at 6%?
---
General Formula: A(t) = P(1+(r/n))^(nt)
-----
Your problem:
2P = P(1+ (0.06/1))^(1*t)
---
(1.06)^t = 2
----
Take the ln of both sides:
t = ln(2)/ln(1.06) = 11.9 years
====================================
Cheers,
Stan H.
===========